OpenStudy (anonymous):

Which of the following points is an equal distance (equidistant) from A(0, −4) and B(−2, 0)? J(−4, −5) K(−3, 0) M(0, 0) N(3, 0)

OpenStudy (danjs):

are you practicing using the distance formula now as the subject? \[d = \sqrt{(x2-x1)^2+(y2-y1)^2}\]

OpenStudy (anonymous):

I used the distance formula and the mid-point formula. And neither of them has given me the answer

OpenStudy (danjs):

k did you do minus a negative as addition wrong maybe, a - (-b) = a+ b

OpenStudy (anonymous):

Yes for the distance formula I got 20. For the mid-point formula I got (-1,-2) You can check my answers

OpenStudy (danjs):

do your distance formula for the last point N(3 , 0) distance from (0, -4) to N and distance from (-2, 0) to N

OpenStudy (anonymous):

OK one sec

OpenStudy (anonymous):

They are both 25. So does that mean it's (3,0)?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

both are root(25) = 5 the distance is the same from both given points to point N

OpenStudy (danjs):



OpenStudy (anonymous):

OK so (3,0) is the equal distant! Thanks!

OpenStudy (danjs):

\[\sqrt{(0-3)^2+(-4-0)^2} = \sqrt{(-2-3)^2+(0-0)^2} = \sqrt{25} = 5\]

OpenStudy (danjs):

yep, not sure where you went wrong to begin with, but you calculated it correct now